# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v1

@@ 3.5 Final Loss Function for Unlearning
In practice, training (unlearning) for the LLM model involves sampling corpora from each dataset, $D_{f}$ and $D_{r}$, and minimizing the final loss function through stochastic gradient descent.
@@ 4 Experiments
In this section, we first perform unlearning experiments based on the Training Data Extraction Challenge (§4.1), followed by ablation and analytical results (§4.2). All experiments were conducted on a remote GCP server equipped with 16 NVIDIA A100 Tensor Core GPUs.
Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset. Note that, in all experiments using LoRA, GD is used as the default loss function. The "+IHL" results refer to experiments where GA within GD is replaced with the proposed IHL (i.e., using Eqn.(5)), and "+FLoRA" indicates the results where FLoRA is used instead of LoRA.
@@ 4.1 Training Data Extraction Challenge
Assuming a scenario where an adversary attacks a pretrained LLM to extract text samples previously used in training, the Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) that are found to be easily extractable. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the validation subset of WikiText as it contains factual world knowledge that we wish to maintain within the LLM.
For TDEC experiments, we consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using three different forget sets. For this experiment, we use a fixed learning rate of 5e-5 when full finetuning, and 2e-4 when using LoRA and FLoRA. For all model sizes, we use a LoRA (and FLoRA) rank of $r=16$ as a default. We set the default to applying LoRA to the attention matrices for Query (Q) and Value (V), as well as to the Feedforward Network (FFN).
@@
Following previous work (Jang et al., 2023), we measure the unlearning efficacy via two metrics. The ${\bm{n}}$-gram Extraction Likelihood (ELn) measures the $n$-gram overlap between the ground truth sequence in ${\mathcal{D}}_{f}$ and the output generated by the model. The Memorization Accuracy (MA) measures the token-wise accuracy of the LLM on ${\mathcal{D}}_{f}$. More details on these metrics are introduced in Section B of the Appendix. After each unlearning epoch, we measure EL10/AM and consider the model to have successfully unlearned ${\mathcal{D}}_{f}$ if both values are smaller than those measured using a held-out validation set. Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to ensure the LLM maintains its previously acquired language modeling capabilities after unlearning. To assess its reasoning capabilities, we average accuracies across 9 different classification datasets LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), PubmedQA (Jin et al., 2019). To measure generative performance, we also measure the F1 score averaged over four dialogue generation datasets, Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues Rashkin et al. (2019), Wizard of Internet (Komeili et al., 2021), Wizard of Wikipedia (Dinan et al., 2018). Lastly, we measure the perplexity on the validation subset of the Pile (Gao et al., 2020).
Baseline.
For experiments on knowledge unlearning in LLMs, we select using two baseline algorithms, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), which rely solely on the original language model while utilizing a forget dataset $D_{f}$ only or both a forget dataset $D_{f}$ and an auxiliary dataset $D_{r}$, respectively. For all experiments, we use the same forgetting criteria for ELn and MA as in  Jang et al. (2023) and report the results after performing unlearning until these criteria are met.
@@
Figure 3: Results from unlearning 32 examples from Training Data Extraction Challenge. Each row represents the average performance in each set of LM capabilies: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). Each column shows results under different model size: 125M, 1.3B, and 2.7B. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines show respective performances prior to unlearning. Unlearning with vanilla GD not only leads to significant loss in performance, but also fails to unlearn sufficiently even with large LoRA ranks. Replacing the negative cross-entropy loss with our hinge loss (IHL) shows significant gains in reasoning and generation capabilities, but still fails to unlearn a number of example sets. Using NCE but with Fisher-weighted LoRA initialization (FLoRA) notably increases the rate of successful unlearning, but at the cost of significant loss in overall LM capability. Using both IHL and FLoRA best minimizes post-unlearning performance degradation in all three capabilities under successful unlearning.
Results.
@@
Table 1 shows the evaluation results of full-parameter unlearning and unlearning with LoRa and FLoRA. From our experimental results, we can identify the following findings. First, for all model sizes, using GD for unlearning not only achieves the forgetting criteria in fewer epochs compared to using GA alone, but it also better preserves previously acquired knowledge. Specifically, for the 125M and 1.3B models, knowledge unlearning with GA leads to a significant deterioration in PPL on the Pile dataset, whereas GD not only overcomes this issue but also achieves improved Dialogue (F1) performance compared to Before (e.g., GPT-Neo 1.3B). Second, we observe that simply applying GD loss with LoRA fails to achieve successful unlearning across all model sizes. Despite the fact that only about 0.7% to 1.6% of the total parameters are updated for unlearning when using LoRA, its simple application actually causes forgetting of previously acquired knowledge. Finally, when using the proposed IHL along with FLoRA, it achieves the forgetting criteria in fewer epochs while better preserving previously acquired knowledge. Although using IHL with LoRA alone maintains previously acquired knowledge better than the baseline, it requires relatively more epochs to meet the forgetting criteria. However, when combined with FLoRA, not only is the number of epochs required to meet the forgetting criteria significantly reduced, but overall, previously acquired knowledge is preserved more successfully.
Additionally, we conduct further experiments with different ranks for LoRA and FLoRA, with results and detailed explanations provided in Figure 3 and its caption. From all experiments so far, we could confirm that simply applying LoRA to GD does not lead to successful unlearning. However, using the proposed IHL in combination with FLoRA demonstrates better cost-efficiency than the existing GA and GD methods, while also better preserving previously acquired knowledge during unlearning.
Figure 4: Results from unlearning 32 examples from Training Data Extraction Challenge using LoRA for GPT-Neo 2.7B. Note that each color denotes Unlearning with vanilla GD, Replacing the negative cross-entropy loss with our hinge loss (IHL), Using NCE but with Fisher-weighted LoRA initialization (FLoRA), and Using both IHL and FLoRA. Each circle represents the result after successfully unlearning 32 examples chosen under a particular random seed. X-marks indicate unsuccessful unlearning attempts based on MA and EL10 metrics.
@@ 4.2 Analysis
Figure 4 presents experiments where LoRA or FLoRA is applied to various weight matrices, including those for Query (Q), Value (V), Key (K), Output (O) in the self-attention module, and the FeedForward Network (FFN). While the original LoRA paper indicates that applying LoRA to Q and V yields superior performance on downstream tasks (Hu et al., 2021), our experiments show that applying LoRA to Q and V alone is insufficient to meet the unlearning criteria in all cases. Notably, applying LoRA to the FFN results in much more successful unlearning in cases of using FLoRA. Furthermore, applying FLoRA to QVFFN in conjunction with IHL not only achieves superior unlearning results but also better preserves previously acquired knowledge. Based on these findings, we adopt the application of LoRA and FLoRA to QVFFN as the default setting.
Cost-efficiency of the proposed method.
@@
To compare unlearning costs, we evaluate the FLOPs for each method and the results are presented in Figure 2. While using vanilla LoRA offers significant advantages in terms of unlearning cost (i.e., FLOPs), as confirmed by our previous experiments, it results in substantial performance losses compared to full-parameter unlearning. In contrast, combining the proposed Inverted Hinge Loss with FLoRA not only achieves the best performance but also leverages the cost advantages of LoRA, demonstrating that this approach enables cost-efficient unlearning.
@@ 5 Concluding Remarks
In this paper, we address limitations of existing unlearning techniques (i.e., Gradient Ascent) by introducing the novel Inverted Hinge Loss (IHL) and leveraging LoRA with a novel initialization (i.e., Fisher-weighted LoRA Initialization (FLoRA)). Our findings reveal that IHL overcomes key drawbacks associated with Gradient Ascent (GA), such as dispersed gradients and unbounded loss, leading to faster and more robust unlearning. Furthermore, our approach of combining FLoRA with IHL not only reduces the cost of unlearning but also enhances the preservation of previously acquired knowledge. Through experiments, we show that our method achieves superior performance in knowledge unlearning compared to the baseline method.
@@ Appendix A Derivative Analysis for the Inverted Hinge Loss Function
For the first case, using the chain rule again, we get:

### v2

@@ 3.5 Final Loss Function for LLM Unlearning
In practice, training (unlearning) for the LLM model is conducted by minimizing Eq.3 through stochastic gradient descent.
@@ 4 Experiments
In this section, we first perform experiments unlearning samples from the Training Data Extraction Challenge (TDEC; §4.1), followed by ablation and analytical results (§4.2). We also conduct experiments on the Task of Fictitious Unlearning (TOFU; §4.3), a benchmark that well-mimics a real-world scenario for LLM unlearning evaluation. For brevity, we present results from additional experiments such as continual unlearning in Appendix D.
Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset. For all runs under LoRA, we use rank 16 and GD is used as the baseline. The "+IHL" results refer to experiments where NCE loss within GD is replaced with our proposed IHL (i.e., Equation 3), and "+FILA" indicates results where FILA is used to initialize LoRA.
@@ 4.1 Training Data Extraction Challenge
Experimental Setup.    The Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the subset of WikiText (Merity et al., 2017) as it contains factual world knowledge that we wish to maintain after unlearning. We consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using five different forget sets. For this experiment, we use a fixed learning rate of 2e-4 and use LoRA adapters with rank $r=\{4,8,16,32\}$. For reasons we illustrate later in §4.2, we choose to apply LoRA on query and value layers in the attention module and two linear layers within feed-forward layers.
@@
Following previous work (Jang et al., 2023), we measure the unlearning efficacy with two metrics. The ${\bm{n}}$-gram Extraction Likelihood (ELn) measures the $n$-gram overlap between the ground truth sequence in ${\mathcal{D}}_{f}$ and the output generated by the model. The Memorization Accuracy (MA) measures the token-wise accuracy of the LLM on ${\mathcal{D}}_{f}$. More details on these metrics are shared in Appendix B. After each unlearning epoch, we measure EL10 and MA of the model, and we consider the model has successfully unlearned ${\mathcal{D}}_{f}$ if both values measured on ${\mathcal{D}}_{f}$ become smaller than those measured from a held-out validation set that the model has never seen before within 20 unlearning epochs. Once unlearning is finished, we evaluate the unlearned model on various downstream benchmarks to measure how well the LLM maintains its previously acquired reasoning and generative capabilities. To assess its reasoning capabilities, we average accuracies across 9 different classification datasets. To measure generative performance, we also average the F1 scores over four dialogue generation datasets. Lastly, we measure the perplexity on the validation subset of the Pile (Gao et al., 2020). A comprehensive list of evaluation datasets can be found in Appendix C.
We consider two LLM unlearning baselines, Gradient Ascent (GA) (Jang et al., 2023) and Gradient Difference (GD) (Liu et al., 2022a; Maini et al., 2024), both of which only require the original language model and datasets ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$ representing knowledge we wish to unlearn and retain, respectively. We exclude methods that require another auxiliary model (Wang et al., 2023; Liu et al., 2024) or the entire training data (Wang et al., 2023; Chen & Yang, 2023) from our baselines.
@@
Figure 3: Results from unlearning examples in the TDEC dataset from the GPT-Neo LLMs with three different sizes. Each row represents the performance averaged across datasets within each set of LLM capability tests: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). The circles and crosses represent successful and unsuccessful attempts, respectively, of unlearning a particular forget set ${\mathcal{D}}_{f}$. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines indicate the base model performance prior to unlearning. Unlearning with GD leads to significant loss in performance and also frequently fails to unlearn effectively even with large LoRA ranks. Replacing the negative cross-entropy (NCE) loss with our hinge loss (IHL) boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases. Using NCE loss with FILA initialization notably increases the rate of unlearning success, but at the cost of losing overall performance. Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
@@
Results.   Table 1 shows evaluation results using a fixed LoRA rank of 16 and Figure 3 shows analogous results using a different LoRA ranks. Our key findings are as follows. First, using GD not only meets the forgetting criteria in fewer epochs across all model sizes but it also preserves previously acquired knowledge (i.e., performance on Reasoning, Dialogue, and Pile) better than GA. Mainly for 125M and 1.3B models, GA causes a significant decline in generative performance, whereas GD partially mitigates this decline and improves both Dialogue F1 scores and the Pile perplexity. Second, we find that simply applying GD with LoRA fails to achieve successful unlearning across all model sizes. Second, we find that imply running GD with LoRA significantly increases the number of epochs required for successful unlearning. While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in significant loss in overall language capability, especially on generative tasks. Lastly, using our proposed IHL in conjunction with FILA achieves the forgetting criteria in fewest epochs while best preserving LLM performance. Although IHL with LoRA alone retains knowledge better than the baseline, it requires more epochs to meet the forgetting criteria. This is resolved when used together with FILA, and the number of required epochs is significantly reduced, reducing catastrophic forggetting of previously acquired knowledge even further.
Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo 2.7B. The marker shapes and colors are used similarly as in Figure 3. Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
@@ 4.2 Analysis
What modules do we need to adapt?    Figure 4 presents experiments where low-rank adapters are attached to various target parameter groups, including those for Query (Q), Value (V), Key (K), Output (O) in the attention module, and the Feed-Forward Network (FFN). While the original LoRA paper (Hu et al., 2022) indicates that applying LoRA to Q and V yields superior performance on downstream tasks, our experiments indicate that using LoRA on Q and V only is insufficient to meet the unlearning criteria within our timeframe of 20 epochs. Notably, when LoRA is applied to FFNs, we observe significant increase in rate of successful unlearning. Furthermore, integrating FILA with IHL achieves the best post-unlearning performance across all LoRA target module combinations.
@@
Cost-efficiency of the proposed method.    Our compute-cost vs. performance comparisons in Figure 2 show that, while vanilla LoRA allows significant reduction in unlearning costs (i.e., FLOPs) by freezing the majority of parameters, it incurs substantial performance losses compared to full-parameter unlearning due to excessive stability originating from its low-rankness. In contrast, combining the proposed IHL with FILA not only achieves the best performance but also leverages the cost advantages of LoRA.
4.3 Task of Fictitious Unlearning
Experimental Setup.  The Task of Fictitious Unlearning (TOFU) benchmark (Maini et al., 2024) is a synthetic dataset containing 20 question-answer pairs for each of 200 fictitious author profiles generated by GPT-4. The TOFU evaluation pipeline first finetunes a pretrained LLM on all QA pairs. Given this finetuned LLM that serves as our base model, our task is to unlearn all information regarding 1%, 5%, or 10% of the authors from the model. Note that we can obtain reference models finetuned only on the retain set (QA-pairs on 99%, 95%, or 90% of authors), with which we evaluate the Forget Quality of unlearned models by measuring the p-value from a Kolmogorov-Smirnov test. A high p-value indicates high distributional similarity between the unlearned model and the reference model, thus implying strong forgetting. To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance. Further details on the dataset and evaluation pipeline can be found in  Maini et al. (2024).
Following the original paper of TOFU, we prepare two base models by finetuning Phi-1.5B and Llama2-7B on TOFU for 5 epochs with learning rates 2e-5 and 1e-5, respectively. We then unlearn using two baselines (GA and GD) and our two methods (IHL and IHL+FILA) using LoRA adapters of rank 4, 8, 16, or 32. For unlearning, we use a learning rate of 2e-4 if our base model is from Phi-1.5B and 1e-4 for Llama2-7B. All training procedures run 5 epochs with an effective batch size of 32 using the AdamW optimizer (Loshchilov & Hutter, 2019).
(a) Phi-1.5B
(b) Llama2-7B
[…]
Results. Figure 5 shows the model utility vs. forget quality curves from unlearning three differently-sized TOFU forget sets from Phi-1.5B and Llama2-7B models. Comparing results among different forget set sizes, we first observe that forgetting 1% of author profiles is fairly straightforward, as all curves quickly approach the reference model with a single epoch, with increasing the LoRA rank leading to incremental improvements in performance. On the other hand, when unlearning a larger set of profiles (i.e., 5% or 10%), we see that both GA and GD quickly degrades model utility.
With regards to our proposed method, we find that replacing the NCE loss in GD with our IHL better retains model utility across all LoRA ranks and forget set sizes, as curves are more aligned straight-up towards the reference point with negligible shift in model utility. This stability comes at the cost of unlearning efficiency, however, as the rate at which the LLM forgets ${\mathcal{D}}_{f}$ is slower with IHL due to IHL decreasing the likelihood of the unwanted token by increasing the likelihood of only one other most-possible token each time in a controlled manner. Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.
Interestingly, we find the prior weight assignment via FILA can lead to excessive unlearning in some cases (e.g., unlearning 10% forget set with ranks 8 or 16 on Llama2-7B), with model updates reducing the forget quality after reaching the upper bound at zero. This behavior resembles the Streisand effect as unlearning gradients beyond a certain point in optimization unintentionally renders ${\mathcal{D}}_{f}$ more noticeable within the model (Golatkar et al., 2020). As reference models are not available for measuring forget quality in real-world scenarios, finding the optimal point at which to stop unlearning to prevent this effect as well as designing a robust evaluation metric that does not depend upon oracle models would be interesting directions, which we leave as future work.
@@ 5 Concluding Remarks
In this paper, we address limitations of Gradient Ascent (GA), a widely used method for LLM unlearning, and introduce a novel Inverted Hinge loss (IHL) to replace the negative cross-entropy loss in GA and resolve issues with dispersed gradients and unboundedness. We also propose Fisher-weighted initialization for low-rank adaptation (FILA) that pre-assigns weights relatively important to generating unwanted information as means to facilitate efficient LLM unlearning with LoRA. Experiments on the Training Data Extraction Challenge dataset with GPT-Neo models along with the TOFU benchmark using Phi-1.5B and Llama2-7B models show that our proposed methods enable faster and more stable LoRA-based LLM unlearning, significantly outperforming existing baselines in computational efficiency as well as post-unlearning performance.
@@ Appendix A Derivative Analysis for the Inverted Hinge Loss Function
For the second case, using the chain rule again, we get:
@@ Appendix B Evaluation Metrics
Appendix C Additional Details on Experimental Setting
Experiemtal Settings   All experiments were conducted on a remote server equipped with NVIDIA A100 40GB Tensor Core GPUs.
Datasets for Evaluation in the TDEC   To evaluate reasoning capabilities, we utilize nine different classification datasets: LAMBADA (Paperno et al., 2016), Hellaswag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), COPA (Gordon et al., 2012), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), PiQA (Bisk et al., 2020), MathQA (Amini et al., 2019), and PubmedQA (Jin et al., 2019). To assess generative performance, we employ Blended Skill Talk (Smith et al., 2020), Empathetic Dialogues (Rashkin et al., 2019), Wizard of Internet (Komeili et al., 2022), and Wizard of Wikipedia (Dinan et al., 2019).
Details of metrics of TOFU   We evaluate the Forget Quality of unlearned models by measuring the p-value from the Kolmogorov-Smirnov test that compares the empirical distribution of our unlearned model to that of the reference model. To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance on the retain set of remaining fictitious author profiles, and two held-out sets consisted of QA-pairs regarding real author profiles and other world facts.
Appendix D Additional Experimental Results
D.1 Continual Unlearning
Because of the importance of continual unlearning (or sequential unlearning) in real-world applications, previous studies have underscored its relevance through a sequence of unlearning tasks (Cha et al., 2024; Jang et al., 2023). Building on them, we conduct continual unlearning experiments involving four tasks. Figure 6 of the Appendix shows that IHL consistently outperforms GD across all metrics. Notably, the proposed IHL demonstrates significantly enhanced performance on the four Dialogue and Pile datasets. Finally, we confirm that the combination of IHL and FLoRA achieves more robust and cose-efficient continual unlearning, as evidenced by the experimental results for Reasoning, Dialogue, and Pile, while utilizing only about 1.6% of the total parameters.
Figure 6: Experimental results of continual unlearning. Each task consists of 32 disjoint sequences sampled from the TDEC dataset, leading to a total of 128 sequences to unlearn. For these experiments, we use the pretrained GPT-Neo 125M model. The experimental setup for unlearning and the forgetting criteria are configured as in the previous TDEC experiments. Task 0 refers to the result before unlearning.